OpenStudy (anonymous):

The numerator and denominator of a fraction are in the ratio of 3 to 5. If the numerator and denominator are both increased by 2, the fraction is now equal to 2/3. If n = the numerator and d = the denominator, which of the following systems of equations could be used to solve the problem? 5n = 3d and 3n + 6 = 2d + 4 5n = 3d and 4n + 4 = 3d + 6 3n = 5d and 3n + 6 = 2d + 4

OpenStudy (anonymous):

they dont becuz i blew ur problem up

OpenStudy (anonymous):

wat u mean

OpenStudy (anonymous):

well u see, u went to ur bathroom and then i snuck into ur room and then blew the pellet outta of u

OpenStudy (anonymous):

whos this r u in my class right now

OpenStudy (anonymous):

hahahahaha

OpenStudy (anonymous):

if so stand

OpenStudy (anonymous):

lol wut

OpenStudy (anonymous):

stand up an come give me a hug

OpenStudy (anonymous):

im not gay unfortunately

OpenStudy (anonymous):

for you

OpenStudy (anonymous):

so ur a guy or girl

OpenStudy (anonymous):

thx @alanli123 is being a creep

OpenStudy (shubhamsrg):

n/d = 3/5 and (n+2)/(d+2) = 2/3 why dont you try and cross multiply and simplify yourself..me and @alanli123 are always there to help in case of any difficulty!! :D

OpenStudy (anonymous):

love u john <3

OpenStudy (anonymous):

i might love u to

OpenStudy (anonymous):

i might just stand up and hug u right now

OpenStudy (shubhamsrg):

no @alanli123 ,,,@johhjohn is mine,,stay away!! ;)

OpenStudy (anonymous):

hey john do u wanna boom chica boom boom with my lica lica HOO HOO

OpenStudy (anonymous):

maybe

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

so wats the answer

OpenStudy (shubhamsrg):

hehehehe..didnt you try yourself? o.O

OpenStudy (anonymous):

yes but failed it singed me out a gave me the wrong answer

OpenStudy (shubhamsrg):

and your ans was?

OpenStudy (anonymous):

b

OpenStudy (saranya):

The equations are: n/d =3/5 and (n+2)/(d+2) = 2/3 If you cross multiply both, you get 5n=3d and 3n + 6 = 2d +4, hence the first one is correct. :)

OpenStudy (shubhamsrg):

there we go!! :D

OpenStudy (anonymous):

haha yay thx